OpenStudy (hotmama2):

What is the slope of the graph of 6x + 12y = 7?   A. -2   B. -1/2   C. 1/2   D.2

OpenStudy (hotmama2):

please help ill fan and metle

OpenStudy (texaschic101):

put your equation in y = mx + b form, and m is your slope

OpenStudy (texaschic101):

6x + 12y = 7-- subtract 6x 12y = -6x + 7 -- divide by 12 y = -6/12x + 7/12 y = -1/2x + 7/12 the number in the m position (the slope) = - 1/2 understand ?